OpenStudy (anonymous):

79.8 is what percent of 95

OpenStudy (anonymous):

79.8 = (x/100) * 95 (x/100) = 79.8/95 (x/100) = 0.84 x = 0.84 *100

OpenStudy (anonymous):

x = 84

OpenStudy (anonymous):

its 84%

OpenStudy (anonymous):

yesssssss